OpenStudy (anonymous):

There is an 80% chance of rain tomorrow and a 20% chance the football game will be postponed. What is the probability that it will rain tomorrow and the football game will be postponed 4% 16% 20% 60%

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

im pretty positive not 100%

OpenStudy (anonymous):

60%

OpenStudy (anonymous):

It cannot be 60% when one of the probability is lower than 60% itself

OpenStudy (anonymous):

BC. 60% is closer to 100

OpenStudy (anonymous):

And they want to know the probability of BOTH occurring.

OpenStudy (anonymous):

is it 20%

OpenStudy (anonymous):

Oh yeah ur right !! Sorry

OpenStudy (anonymous):

@shalante

OpenStudy (anonymous):

i read it wrong. dont listen to me XD

OpenStudy (anonymous):

Since 80% is 4/5 and 20% is 1/5 \[\frac{ 4 }{ 5 }*\frac{ 1 }{ 5 }=\frac{ 4 }{ }\]

OpenStudy (anonymous):

oh wow that denominator is bugged again. supposed to be 4/25

OpenStudy (anonymous):

16

OpenStudy (anonymous):

or just multiply the percentage 80%=0.80 20%=0.20 (0.80)(0.20)

OpenStudy (anonymous):

yea 16% (0.16) would equal 16%

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem!